OpenStudy (anonymous):

What is the slope and y- intercept of the graph of 2x+3y=18?

OpenStudy (bahrom7893):

3y = -2x + 18 y = (-2/3)x + 6 Slope = -2/3 Y intercept: (0,6)

OpenStudy (anonymous):

^ bah is right :)